OpenStudy (anonymous):

Please help!!!!!!!!!!!!!!!! Give 3 or 4 examples of simplifying Radicals. Explain and show work. Also what does simplifying radicals mean?

OpenStudy (anonymous):

Any help? :)

OpenStudy (anonymous):

i think it be best if u give the examples and we help u solve them....

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I don't have any examples, I need to give at least 3 examples and solve them. It can be any examples you can think of

OpenStudy (anonymous):

simplifying radicals means that the radicand does not have any perfect squares as one of its factors.

OpenStudy (anonymous):

ok.... i'll give you one example.... simplify: \(\Large \sqrt{56} \) \(\Large \sqrt{56}=\sqrt{4\cdot14}=\sqrt{4}\cdot \sqrt{14}=2\sqrt{14} \)

OpenStudy (anonymous):

in this example, since 4 is a perfect square number and is a factor of the radicand, it can be pulled out of the radical symbol... now the simplified form does not have a perfect square as a factor in the radicand.

OpenStudy (anonymous):

Ok thanks . Any more examples it does have ti be long and hard it can be simple and easy

OpenStudy (anonymous):

What is a radicand?

OpenStudy (anonymous):

the radicand is the number INSIDE the radical symbol... so in my example, 56 is the radicand

OpenStudy (anonymous):

Do you have another example?

OpenStudy (anonymous):

why don't u give me any number and we'll simplify the square root (or any root of) of that number....

OpenStudy (anonymous):

Okay how about the square root of 3?

OpenStudy (anonymous):

ok... \(\Large \sqrt3 \) the radicand is 3... since 3 is already prime (it's factors are 1 and itself), \(\sqrt3 \) is already in it's simplified form

OpenStudy (anonymous):

\(\Large \sqrt3 + \sqrt2 \) is also already written in simplified form... if you want an APPROXIMATE value for this sum, use a calculator.

OpenStudy (anonymous):

Can you simplify the square root of 28? And the square root of 132